### v5

# A Unified Framework for Automated Code Transformation and Pragma Insertion

Abstract.
High-Level Synthesis compilers and Design Space Exploration tools have greatly advanced the automation of hardware design, improving development time and performance. However, achieving a good Quality of Results still requires extensive manual code transformations, pragma insertion, and tile size selection, which are typically handled separately. The design space is too large to be fully explored by this fragmented approach. It is too difficult to navigate this way, limits the exploration of potential optimizations, and complicates the design generation process.
To tackle this obstacle, we propose Sisyphus, a unified framework that automates code transformation, pragma insertion, and tile size selection within a common optimization framework. By leveraging Nonlinear Programming, our approach efficiently explores the vast design space of regular loop-based kernels, automatically selecting loop transformations and pragmas that minimize latency.
Evaluation against state-of-the-art frameworks, including AutoDSE, NLP-DSE, and ScaleHLS, shows that Sisyphus achieves superior Quality of Results, outperforming alternatives across multiple benchmarks. By integrating code transformation and pragma insertion into a unified model, Sisyphus significantly reduces design generation complexity and improves performance for FPGA-based systems.
HLS, code transformation, pragma insertion, non-linear problem
journalyear: 2025copyright: rightsretainedconference: Proceedings of the 2025 ACM/SIGDA International Symposium on Field Programmable Gate Arrays; February 27–March 1, 2025; Monterey, CA, USAbooktitle: Proceedings of the 2025 ACM/SIGDA International Symposium on Field Programmable Gate Arrays (FPGA ’25), February 27–March 1, 2025, Monterey, CA, USAdoi: 10.1145/3706628.3708873isbn: 979-8-4007-1396-5/25/02ccs: Hardware High-level and register-transfer level synthesisccs: Software and its engineering Compilers
1. Introduction
High-Level Synthesis (HLS) compilers and source-to-source compilers are indispensable tools in accelerating hardware design by automating the translation of high-level programming languages like C/C++ or Python into hardware descriptions. They offer notable advantages, such as reduced development time and enhanced performance for hardware designs (Cong et al., 2011; Zhang et al., 2008; Xilinx, 0232; Intel, 2024; Siemens, 2023; Microchip, 2023; Xilinx, 2024; Huang et al., 2021; Ye et al., 2022; Lai et al., 2019; Yu et al., 2018; Li et al., 2020; Zhang et al., 2024; Cong et al., 2022). However, achieving a good quality of results (QoR) often requires manual code transformations and pragma insertions, which can be facilitated with either Design Space Exploration (DSE) (Pouget et al., 2024, 2025; Sohrabizadeh et al., 2022b, 2023, a; Ferretti et al., 2018; Zhao et al., 2017; Sohrabizadeh et al., 2023) or source-to-source compilers (Lai et al., 2019; Ye et al., 2022, 2024; Zhang et al., 2024) to guide the synthesis process. These pragmas aid in optimizing the generated hardware code.
Despite their advantages, existing HLS workflows face several significant challenges. Previous frameworks (Ye et al., 2022, 2024; Lai et al., 2019; Zhang et al., 2024) have integrated pragma insertion and code transformation, but these efforts are somewhat limited, focusing primarily on loop permutations based on loop properties. Additionally, code transformations and pragma insertion are often treated as separate processes. Transformations such as loop tiling, splitting, and permutation are typically performed in advance or as part of a pre-processing stage, followed by pragma insertion for optimization, like loop unrolling, pipelining and array partitioning. This separation introduces inefficiencies, as transformations can affect the effectiveness of pragmas, and vice versa. Moreover, the immense design space—encompassing millions of potential transformation and pragma combinations—makes exhaustive exploration unfeasible. Ensuring the correctness of these transformations while adhering to resource constraints further complicates the process.
Our primary research objective is to develop a system capable of autonomously conducting code transformation, tile size selection to cache the data on-chip and integrating hardware pragmas for HLS to enhance parallelization. We seek to attain a favorable QoR, especially for computation-bound designs, where maximizing parallelism is crucial for optimizing QoR. To address this challenge, we introduce Sisyphus, a framework built on top of HLS compilers. This framework automates code transformation, including loop splitting, permutation, tiling, and pragma insertion for unrolling, pipelining, and on-chip data caching while partitioning arrays to ensure efficient parallelization, all within a single optimization framework. Even though several exploration methods could be used, we chose to employ a Nonlinear Programming (NLP) approach. This method facilitates rapid exploration of the entire theoretical space. We have developed an analytical model that integrates considerations of latency and resource utilization, building upon previous research (Pouget et al., 2025, 2024). Our focus lies specifically on regular loop-based kernels (PolyBench, [n. d.]), ensuring meticulous control over the correctness of code transformations and the accuracy of cost models. This model relies on parameters derived from both the program’s schedule and the pragma configuration. To facilitate seamless code transformation while respecting constraints and pragma insertion, we have designed a novel template tailored to these objectives. In our template, we incorporate a two-level tiling strategy, where each tiling level corresponds to particular HLS optimizations such as fine-grained unrolling, pipelining, coarse-grained unrolling, or tiling. As a result, the loop trip counts become variables, with each loop associated with a specific pragma determined by the tiling level. By solving the NLP problem, we determine these trip counts, as well as other parameters like the array partitioning factor, enabling the automatic generation of the corresponding C++ code.
In summary, we introduce the following contributions:
  • A novel optimization template that integrates code transformation, pragma insertion, and tiles size selection for data caching by consolidating these tasks into a single optimization problem. This approach not only simplifies the search space but also ensures that only legal transformations are considered.
  • A novel NLP-based approach specifically designed to explore the joint design space of regular loop-based kernels. Unlike traditional methods that involve time-consuming DSE, our approach provides solutions within minutes. Plus, the NLP is user-friendly, allowing users to easily configure parameters, such as on-chip buffer size, to meet the specific requirements of the kernel.
  • Our framework acts as a comprehensive, fully automated system, offering end-to-end functionality. With it, we can conduct thorough evaluations and attain QoR designs that are comparable or superior to those achieved by AutoDSE, NLP-DSE, ScaleHLS and HARP – all without requiring DSE to perform multiple HLS compiler syntheses. Furthermore, Sisyphus identifies the correct transformation, even when no transformation is needed.
The paper is organized into the following sections. Section 2 explains our approach and solution. Following that, Section 3 covers code transformation and pragma insertion. Section 4 introduces a nonlinear formulation based on this model to automatically discover schedule and pragma configurations through NLP optimization. Section 5 elaborates on our post-optimizations. Finally, sections 6, 7, 8 and 9 are dedicated to evaluating our method, presenting related work, addressing limitations and drawing conclusions.
2. Background and Motivation
2.1. Design Space to Explore
The exploration space for optimizing an HLS design includes pragma insertion, code transformations, and data caching, all within the limits of resource constraints. Although different objective functions can be explored, this work specifically focuses on minimizing latency while respecting resource constraints. This objective function is especially potent for computation bound kernels because it addresses the bottleneck caused by insufficient parallelization.
The various components of the exploration space—pragma insertion, code transformations, and data caching—are inherently interdependent. For example, unrolling pragmas affect array partitioning, which in turn influences BRAM utilization. Similarly, loop scheduling and pragmas impact the potential for on-chip data caching, while the choice of schedule dictates how loops should be unrolled based on their characteristics, such as whether they are reduction loops. A reduction loop is a loop where an operation, like addition or multiplication, accumulates values across iterations to produce a single output, often by reducing a multi-element array to a scalar (e.g., loop L2 in Listing 1). The choice of schedule for reduction loops, as opposed to non-reduction loops (e.g., L0, L1, and L3 in Listing 1), influences unrolling strategies and overall optimization because of the dependencies inherent to the loop’s characteristics.
Selecting one optimization affects others, often requiring backtracking when spaces are handled separately. This interdependence complicates exploration and demands constant verification of transformation legality.
1for (i = 0; i < 200; i++) { // L0
2 for (j = 0; j < 220; j++) // L1
3 C[i][j] *= beta; // S0
4 for (k = 0; k < 240; k++) // L2
5 for (j = 0; j < 220; j++)// L3
[…]
9for (i = 0; i < 200; i++) // L0
10 for (j = 0; j < 220; j++) // L1
11 C[i][j] *= beta; // S0
12for (i = 0; i < 200; i++) // L2
13 for (k = 0; k < 240; k++) // L3
14 for (j = 0; j < 220; j++) // L4
[…]
18for (i0 = 0; i0 < I0_S0; i0++) // L0
19 for (i1 = 0; i1 < I1_S0; i1++) // L1
20 for (i2 = 0; i2 < I2_S0; i2++) // L2
21 for (j0 = 0; j0 < J0_S0; j0++) // L3
22 for (j1 = 0; j1 < J1_S0; j1++) // L4
23 for (j2 = 0; j2 < J2_S0; j2++) // L5
24 C[i][j] *= beta; // S0
25for (i = 0; i < I0_S1; i++) // L6
26 for (i = 0; i < I1_S1; i++) // L7
27 for (i = 0; i < I2_S1; i++) // L8
28 for (k = 0; k < K0_S1; k++) // L9
29 for (k = 0; k < K1_S1; k++) // L10
30 for (k = 0; k < K2_S1; k++) // L11
31 for (j = 0; j < J0_S1; j++) // L12
32 for (j = 0; j < J1_S1; j++) // L13
33 for (j = 0; j < J2_S1; j++) // L14
[…]
38
[…]
42On-chip mem = 320 kB}]
43load_C(C, C_off_chip);
44/**** Level 0 of S0 ****/
45for (j0 = 0; j0 < 1; j0++) // L0
46 for (i0 = 0; i0 < 4; i0++) // L1
47/**** Level 1 of S0 ****/
48 for (j1 = 0; j1 < 22; j1++) // L2
49#pragma HLS pipeline
50/**** Level 2 of S0 ****/
51 for (j2=0; j2<10; j2++) { // L3
52#pragma HLS unroll
53 for (i2=0; i2<50; i2++) // L4
54#pragma HLS unroll
55 i = ...;
56 C[i][j] *=beta; } // S0
57/**** Level 0 of S1 ****/
58for (k0 = 0; k0 < 48; k0++) // L5
59 load_B_S1(B, B_off_chip, k0);
60 load_A_S1(A, A_off_chip, k0);
61 for (j0 = 0; j0 < 1; j0++) // L6
62 for (i0 = 0; i0 < 4; i0++) // L7
63/**** Level 1 of S1 ****/
64 for (j1=0; j1 < 220; j1++) // L8
65#pragma HLS pipeline
66/**** Level 2 of S1 ****/
67 for (j2=0; j2<1; j2++) // L9
68#pragma HLS unroll
69 for (i2=0; i2<50; i2++) // L10
70#pragma HLS unroll
71 for(k2=0;k2<5;k2++){ // L11
72#pragma HLS unroll
73 i = ...;
74 C[i][j]+=alpha*A[i][k2]*B[k2][j];}// S1
75store_C(C, C_off_chip);
Listing 1: Original gemm Code
1load_C(C, C_off_chip);
2load_A(A, a_off_chip);
3load_B(B, B_off_chip);
4/**** Level 0 of S0 ****/
5for (i0 = 0; i0 < 1; i0++) // L0
6 for (j0 = 0; j0 < 1; j0++) // L1
7/**** Level 1 of S0 ****/
8 for (j1 = 0; j1 < 55; j1++) // L2
9#pragma HLS pipeline
10/**** Level 2 of S0 ****/
11 for (i2 = 0; i2 < 200; i2++) // L3
12#pragma HLS unroll
13 for (j2=0; j2<4; j2++) { // L4
14#pragma HLS unroll
15 i = ...;
16 C[i][j] *= beta;} // S0
17/**** Level 0 of S1 ****/
18for (i0 = 0; i0 < 1; i0++) // L5
19 for (j0 = 0; j0 < 1; j0++) // L6
20 for (k0 = 0; k0 < 60; k0++) // L7
21/**** Level 1 of S1 ****/
22 for (j1 = 0; j1 < 220; j1++) // L8
23#pragma HLS pipeline
24/**** Level 2 of S1 ****/
25 for (i2=0;i2<200;i2++) // L9
26#pragma HLS unroll
27 for (j2=0; j2<1; j2++) // L10
28#pragma HLS unroll
29 for(k2=0;k2<4;k2++){ // L11
30#pragma HLS unroll
31 i = ...;
32 C[i][j]+=alpha*A[i][k]*B[k][j];}//S1
33store_C(C, C_off_chip);
Listing 2: Result found for gemm kernel using the NLP with the constraints: DSP = 6,840 On-chip mem = 7.2 MB
2.2. Limitation of the Current DSE Methods
Several frameworks (Lai et al., 2019; Ye et al., 2022; Zhao et al., 2022; Ye et al., 2024; Pouchet et al., 2013a; Zhang et al., 2024; Chen et al., 2024) enable code transformations and pragma insertion, but have limited exploration spaces (Ye et al., 2022; Zhao et al., 2022; Ye et al., 2024; Pouchet et al., 2013a; Zhang et al., 2024) or lack automatic DSE (Lai et al., 2019; Chen et al., 2024). Transformations are often restricted to loop property-based permutations (Lai et al., 2019; Ye et al., 2022, 2024; Pouchet et al., 2013a; Zhang et al., 2024) or rely on Pluto (Bondhugula et al., 2008), which is optimized for CPUs but overlooks FPGA-specific techniques like pipelining and array partitioning. Additionally, Pluto minimizes memory transfer from off-chip to on-chip communication (Comm), which can limit parallelism on FPGAs, unlike our focus on reducing global latency (Lat). Meanwhile, other DSE frameworks (Sohrabizadeh et al., 2022b, 2023, a; Pouget et al., 2025) explore pragmas insertion for fixed loop orders, but predicting friendly transformations beforehand is challenging. Exploring every possible transformation is also impractical for large design spaces. The key differences between the frameworks are summarized in Table 1.
Poly-Opt-HLS POL-SCA/ Pluto Auto-DSE/ HARP Hetero-CL / Allo NLP-DSE Scale-HLS/ POM Sisy-phus
Tiling Limit.
Permutation Limit. Limit. Limit.
Pragma Insert.
Code Trans + Pragma Insert. (unified)
Off-chip comm. gen.
Enum. (AI, heuristics,…) Manual
Objective Comm Comm Lat - Lat Lat Lat
Table 1. Comparison of different frameworks
We will now examine examples that highlight the limitations of current methods. For this, we use two different HLS DSE methods, AutoDSE (Sohrabizadeh et al., 2022b) and NLP-DSE (Pouget et al., 2025, 2024). A comparison with HARP (Sohrabizadeh et al., 2023) and ScaleHLS (Ye et al., 2022) will be presented in Section 6. AutoDSE treats the source-to-source compiler Merlin (Xilinx, 2024) and the associated HLS tools as a black box, adjusting pragmas based on the identified bottlenecks from previous iterations. In contrast, NLP-DSE employs Nonlinear Programming DSE, utilizing a lower bound-based objective function to achieve high QoR within a short timeframe. Both of this DSE use the AMD source-to-source compiler Merlin (Xilinx, 2024). The compiler employs different code transformations such as strip-mining via the TILE pragma or for partial loop unrolling. It typically avoids permutations, except when partially unrolling the two innermost loops, in which case the compiler strip-mines these loops and applies permutations to the resulting fully unrolled ones. A more detailed description of this compiler can be found in (Pouget et al., 2025). In the following examples, we use the HLS compiler Vitis 2023.2 with the unsafe-math option disabled. The targeted FPGA is the AMD/Xilinx Alveo U200. We evaluate two scenarios: the gemm kernel from Polybench (PolyBench, [n. d.]), as shown in Listing 1, and a convolutional neural network (CNN) layer where the problem size and original loop order of CNN are I,J=256 (Output and Input Channels), H,W=224 (Height, Weight) and P,Q=5 (Filter Height and Weight). The gemm kernel’s original loop order, as illustrated in Listing 1, has a non-reduction loop at the center of the second statement (L3). This arrangement facilitates AutoDSE and NLP-DSE in unrolling this loop and pipelining the reduction loop (L2). In this scenario, if the reduction loop is pipelined and partially unrolled by a factor of $uf$, performance will degrade. While partial unrolling allows for parallel execution of the multiplications, it increases the pipeline depth due to the loop’s dependencies, as the $uf$ additions must still be performed sequentially. Consequently, when the reduction loop is pipelined, the initiation interval (II) will significantly increase, becoming a multiple of both $uf$ and the latency of the reduction operation. This occurs because the next iteration of the loop can only begin after the $uf$ additions are completed and the output is fully written. Their inability to further increase parallelism due to loop order limits the best designs to a throughput of 20 GFLOPs per second (GF/s). In such instances, a transformation becomes imperative to augment parallelization. A single CNN layer with 6 loops results in an enormous design space that cannot be explored exhaustively. Loop permutations alone represent 6! = 720 possibilities, and tiling is required to manage array sizes exceeding on-chip memory capacity. While frameworks like AutoDSE and NLP-DSE achieve satisfactory throughput—42.15 GF/s and 31.80 GF/s, respectively—they are unable to explore the full design space, which includes both code transformations and pragma insertion, within a reasonable timeframe. For example, running AutoDSE for all possible loop permutations would take around 500 days, and for NLP-DSE, it would take 106 days. Other frameworks, which apply code transformations based solely on loop properties, such as prioritizing reduction loops in the outermost positions, explore a much more restricted design space, limiting their potential for further optimization. This underscores the challenge of navigating such a vast design space and highlights gaps in current approaches.
2.3. Overview of Sisyphus
We introduce Sisyphus, tailored for affine programs with loop bounds and conditionals as affine functions (Girbal et al., 2006; Feautrier, 1988). It defines a unified search space for parameters like pragma insertion, loop ordering, and tiling, avoiding backtracking or disjoint optimizations. Sisyphus efficiently explores this design space, overcoming limitations of frameworks that either overly restrict or partially explore it. Crucially, it ensures only valid transformations. The process involves creating a kernel-specific space and exploring it. The first step involves creating a template with fully distributed code that includes three levels of loops, each targeting specific optimizations. To achieve this, we strip-mined each original loop twice, ensuring that one of these loops is included at each level (assuming the permutation is valid). As a result, each original loop can be utilized across all three levels. For instance, in Listing 1, the loop $k$ (L2) is used in level 0 (L5) and in level 2 (L11) in Listing LABEL:lst:small_mem. Specifically, we have an innermost level for fine-grained unrolling, a level dedicated to pipelining, and another level focused on on-chip data caching based on available resources. As demonstrated in Listings LABEL:lst:small_mem and 2, depending on the resource constraints specified by the user, we obtain two different code versions. For instance, a reduced on-chip memory requirement necessitates the partial transfer of arrays $A$ and $B$ (lines 17, 18 in Listing LABEL:lst:small_mem), leading to the use of a smaller on-chip buffer size that meets the user’s constraints. Once this template is generated, it establishes a search space where the unknowns consist of the loop bounds, loop ordering, buffer sizes for on-chip arrays, and the locations in the code where data transfers take place. Although other DSE methods can explore this space, we chose to use a cost model formulated as a Nonlinear Programming (NLP) problem as in NLP-DSE (Pouget et al., 2025). This approach allows us to efficiently navigate the theoretical space within seconds or minutes, leveraging accurate modeling made possible by compile-time analysis for affine programs. It is important to emphasize that exhaustive exploration using a simple cost model or synthesizing each design with HLS (Sohrabizadeh et al., 2022b) would be impractical due to the sheer size of the search space. In contrast, the NLP approach enables us to explore this space quickly and efficiently, allowing for a more comprehensive investigation than traditional methods. Sisyphus outperforms NLP-DSE (Pouget et al., 2025, 2024) by exploring a larger optimization space, including tiling and loop permutations, unlocking performance unattainable by NLP-DSE, which requires separate runs for each transformation. Automating the entire process, Sisyphus reduces exploration time and delivers superior results. It generates Vitis HLS pragmas with broader parameter support, like array partitioning, enhancing accuracy and design space coverage. Its high model precision enables the generation of at most one or two optimized designs, avoiding NLP-DSE’s extensive iterations. For gemm, Sisyphus significantly enhances parallelism by splitting and permuting loops without relying on the reduction loop. In this configuration, the NLP identifies the setup shown in Listing 2. Both pipelined loops achieve an initiation interval (II) of 1, resulting in a throughput of 210 GFLOPs per seconds (GF/s). The reduction loop (L11) is partially unrolled, allowing the multiplications to be executed in parallel while performing four additions sequentially. This small unrolling, selected by the NLP, deepens the pipeline, improving the quality of results (QoR) even though the reduction must remain sequential. For the CNN, the NLP selects all parameters in the design space, including loop order (exploring 6! possibilities), on-chip array size, parallelism levels, etc. This analysis completes in 13 minutes, achieving a throughput of 341 GF/s. Thanks to a well-defined space that considers only legal transformations and exploration guided by a cost model formulated as an NLP, our framework can identify the theoretical optimum across the entire space and generate a design within seconds or minutes.
3. Unified Space
We proceed to develop a methodology that integrates code transformation, tile size selection and pragma insertion into a unified optimization framework. This involves implementing maximal distribution for code transformation, followed by the design and implementation of a template capable of executing code transformation, tiles size selection and pragma insertion simultaneously. We illustrate our code transformation process step by step using the example from Listings 1 to 2. Input  Sisyphus takes an affine C/C++ file as input and extracts a polyhedral representation using PoCC (PoCC, [n. d.]), gathering information like loop order, bounds, and array sizes. This data is used to create the template and generate the NLP. The framework is most effective for fully distributable and permutable codes but still works with restricted solution space when transformations are not fully legal. Maximal Distribution  To initiate the process, we distribute the program to maximize the distance between statements, promoting parallelization by reducing dependency overlap. Typically, this yields one statement per loop body within perfectly nested loops. Using ISCC (Verdoolaege, 2011), we explore scheduling options and verify transformation legality by preserving dependencies. Listing LABEL:lst:distributed shows the fully distributed version of the code from Listing 1. Strip-mining  After maximal distribution, we apply strip-mining to each loop twice. This process transforms the original loop, such as the loop L0 in Listing LABEL:lst:distributed, into three loops with trip counts $I0\_S0$, $I1\_S0$, and $I2\_S0$ (L0, L1, L2 in Listing LABEL:lst:strip). The product of these three loop trip counts equals the original trip count, i.e., $I0\_S0\times I1\_S0\times I2\_S0=200$. Since strip mining is inherently legal, there is no necessity to validate the legality of this transformation. Subsequently, we can arrange the loops in different permutations if it is legal. After applying three levels of strip-mining, the code from Listing LABEL:lst:distributed is transformed into the code seen in Listing LABEL:lst:strip. Creation of different levels  Next, we consider possible permutations of the strip-mined loops. If these permutations are legal, we establish three loop levels. We ensure legality by checking dependency preservation using ISCC (Verdoolaege, 2011). Level 2:  The innermost level facilitates fine-grained unrolling (complete unrolling), which increases parallelism by duplicating statements and can utilize a tree reduction if a reduction loop is unrolled and the option is enabled. For example, this is illustrated by loops L9, L10, and L11 in Listing 2. Level 1:  The middle level facilitates pipelining, enhancing throughput by overlapping loop iterations. In cases where loop permutations allow us to achieve perfectly nested loops suitable for pipelining, we restrict pipelining to a single loop to ensure efficient synthesis. Flattening the loops at Level 1 into a single loop and then pipelining the resulting loop introduces significant design complexity and leads to excessively long synthesis times. Thus, if a loop at Level 1 is pipelined, it implies that the trip counts of the other loops at the same level are reduced to one. For instance, in Listing 2, the loop $j$ (L8) is pipelined for the statement S1. This results in the loops $k$ and $i$ (L10 and L11 in Listing 3 ) having a trip count of 1, enabling them to be eliminated from the code. Level 0:  Finally, the outermost level manages coarse-grained unrolling, tiling, and/or sequential execution. It not only controls parallelism at a coarse level but also determines the on-chip buffer size through tiling. As described in the following paragraph, the loop order at this level is included in the design space. This level is illustrated by loops L5, L6, and L7 in Listing LABEL:lst:small_mem. The loop bounds determine the on-chip buffer sizes for arrays A and B, transferred below loop L5. Resource constraints also influence the sequential execution of loop L7. If the loop body contains a single statement and loop permutation is feasible, we disable coarse-grained unrolling. This achieves equivalent results since our template allows fine-grained unrolling of the innermost loop. Loop Permutation  Loop order selection is not required for levels 1 and 2, as only a single loop is pipelined, and the innermost level (level 2) is fully unrolled, making the loop order irrelevant. Conversely, at the outermost level (level 0), all permutations are considered, and the NLP selects the permutation. In Listings LABEL:lst:small_mem and 2, we can see that the NLP selected two different permutations for the statement S1 (loops L5, L6, and L7). Fusion  We have opted not to incorporate fusion into our model for several reasons. Firstly, our objective is to maximize parallelization for computation-bound kernels. Fusion may reduce or minimize the dependency distances between statements, limiting potential parallelization opportunities. Additionally, since our focus is on computation-bound kernels, the primary bottleneck lies in computation. Therefore, we are willing to incur a minor cost, even if it involves transferring a tiled array multiple times. Overal Template  In the context of our example, these transformations yield the code presented in Listing 3, with each level corresponding to specific transformations. The iterator of the loop retains its original name, with an added number indicating the level. For example, loop L14 in Listing 3 refers to loop L2 in Listing 1 at level 2 of the strip mining. Notably, our framework is capable of preserving the original loop order, allowing it to retain the existing loop structure if the code is already optimized. Design Space  The challenge is to determine loop trip counts (e.g., $J0\_S0$) and level 0 loop order while meeting DSP and on-chip memory constraints. Hence, the trip counts corresponding to fully unrolled loops must be constrained to avoid over-utilization of DSP resources. The on-chip buffers size must not exceed the on-chip memory capacity. Consequently, for a large problem size, the outermost level will enable control over the buffer sizes.
1int perm_S0[2] // Permutation order (0 for i0, 1 for j0)
2int perm_S1[3] //Array to store permutation order (i0,j0,k0)
3int lb_S0[2] = {I0_S0, J0_S0}; // loop bounds for S0
4int lb_S1[3] = {I0_S1, J0_S1, K0_S1}; // loop bounds for S1
5// possibilities to cache on-chip A, B and C
6/********* Level 0 of S0 *********/
7for (L0_S0=0; L0_S0 < lb_S0[perm_S0[0]]; L0_S0++) // L0
8// possibilities to cache on-chip C
9 for (L1_S0=0; L1_S0 < lb_S0[perm_S0[1]]; L1_S0++) // L1
10 // possibilities to cache on-chip C
11/********* Level 1 of S0 *********/
12 for (j1=0; j1 < J1_S0; j1++) // L2
13 for (i1=0; i1 < I1_S0; i1++) // L3
14#pragma HLS pipeline
15/********* Level 2 of S0 *********/
16 for (i2=0; i2 < I2_S0; i2++) // L4
17#pragma HLS unroll
18 for (j2=0; j2 < J2_S0; j2++) // L5
19#pragma HLS unroll
20 C[i][j]*=beta; // S0
21/********* Level 0 of S1 *********/
22for (L0_S1=0; L0_S1 < lb_S1[perm_S1[0]]; L0_S1++) // L6
23// possibilities to cache on-chip A, B and C
24 for (L1_S1=0; L1_S1 < lb_S1[perm_S1[1]]; L1_S1++) // L7
25 // possibilities to cache on-chip A, B and C
26 for (L2_S1=0; L2_S1 < lb_S1[perm_S1[2]]; L2_S1++) // L8
27 // possibilities to cache on-chip A, B and C
28/********* Level 1 of S1 *********/
29 for (j1=0; j1 < J1_S1; j1++) // L9
30 for (k1=0; k1 < K1_S1; k1++) // L10
31 for (i1=0; i1 < I1_S1; i1++) // L11
32#pragma HLS pipeline
33/********* Level 2 of S1 *********/
34 for (i2=0; i2 < I2_S1; i2++) // L12
35#pragma HLS unroll
36 for (j2=0; j2 < J2_S1; j2++) // L13
37#pragma HLS unroll
38 for (k2=0; k2 < K2_S1; k2++) // L14
39#pragma HLS unroll
40 i = i0 * I1_S1 * I2_S1 + i1 * I2_S1 + i2;
41 j = j0 * J1_S1 * J2_S1 + j1 * J2_S1 + j2;
42 k = k0 * K1_S1 * K2_S1 + k1 * K2_S1 + k2;
[…]
Listing 3: Template automatically generated for the gemm input code (Listing 1)
4. NLP
We automatically generate constraints and variables for a nonlinear program to explore the theoretical solution space outlined in Section 3, adapting the approach from (Pouget et al., 2025, 2024) to our context. Latency is estimated with optimistic DSP utilization, assuming perfect resource reuse, but users can adjust the DSP limit or choose a pessimistic scenario with no reuse. Users can also customize on-chip memory size, maximal array partitioning, and latency/resources per operation, ensuring adaptability to different platforms and compilers. To gather all the required information for the NLP, we utilize PoCC (PoCC, [n. d.]) to extract the intermediate representation.
4.1. Variables
Table 2 defines the sets, variables, and constants utilized in our NLP formulation.
4.2. Constraints
We describe each constraint’s meaning and implications with examples from the paper’s listings. Trip Count  Equation 1 constrains the trip count of each loop, ensuring that the product of the trip counts equals the original trip count. The three loops after strip-mining, e.g., L6, L7, and L8 in Listing LABEL:lst:strip, must be equivalent to the original loop L2 in our example from Listing LABEL:lst:distributed. Therefore, we require $I0\_S0\times I1\_S0\times I2\_S0=200$.
(1) $\forall l\in\mathcal{L},\prod_{v\in\mathcal{V}}TC_{l,v}=TC_{l}$
Set Description
$\mathcal{L},\mathcal{A},\mathcal{S}$ the set of loops, arrays and statements
$\mathcal{S}_{1}$ the set of statements alone in a loop body
$\mathcal{O}_{s}$ the list of operations of the statements $s$
$\mathcal{L}_{s}$ the set of nested loops which iterate the statement $s$
$\mathcal{L}_{s}^{red}$ the set of reduction loops iterating the statement $s$
$\mathcal{C}_{a_{d}}$ the set of loops iterating the array $a$ at dimension $d$
$\mathcal{V}$ Strip mining level: 0 for coarse-grained/sequential, 1 for pipeline, 2 for fine-grained unrolled
$AP_{a,d}$ Array Partition for the array $a$ in dimension $d$
Constant Description
$TC_{l}$ Trip Count of the loop $l$ before strip-mining
$II_{l}$ II of the loop $l$
$IL_{par},IL_{red}$ Iteration Latency of the operations without ($par$) and with ($red$) dependencies of the statement $s$
$DSP_{s_{op}}$ Number of DSP used for the statement $s$ for the operation $op$ (This number accounts for the number of times the statement is replicated due to unrolling)
$DSP_{available}$ Number of DSP available for the FPGA used
$max_{part}$ Maximum array partitioning defined by the user
$ft\_arr_{a}\_loop_{l}$ Footprint of the array $a$ if transferred to on-chip after the loop $l$
$reuse_{opt}$ Boolean for optimistic reuse
Variable Description
$tc_{l,level}$ TC of the loop $l$ for the level of strip-mining
$loop_{l}\_UF$ Coarse-grained unroll factor of the loop $l$ at level 0
$loop_{l}\_pip$ Boolean to know if the loop $l$ is pipelined at level 1
$cache_{l}\_arr_{a}$ Boolean to know if the array $a$ is transferred on-chip after the loop $l$ at level 0
$perm_{s}$ ID of the permutation of the statement $s$
$burst_{a}$ The maximum bitwidth at which the array $a$ can be transferred from off-chip to on-chip
Table 2. Overview of the sets, constants and variables employed in formulating the NLP
Pipeline  Constraints 25 facilitate the selection of loop pipelining. Constraint 2 ensures each loop has a flag indicating if it is pipelined. For example, in Listings LABEL:lst:small_mem and 2, $S1$ shows loop $j1$ (L8) is pipelined, while its original loop is $j$ (L3 in Listing 1), making $loop_{j}\_pip=1$. As per Section 3, only one loop at level 1 can be pipelined, requiring the sum of pipeline flags must be less than or equal to 1 (Eq. 4). Non-pipelined loops must have a trip count of 1 (Eq. 3). For instance, in Listing 3, only loop L9 is pipelined at level 1 of $S1$, so $loop_{j}\_pip=1$, while $loop_{i}\_pip=0$ and $loop_{k}\_pip=0$. Loops L10 and L11 have trip counts of 1 and can be removed. Eq. 5 calculates the initiation interval (II) based on loop properties and reduction operation latency. In Listings LABEL:lst:small_mem and 2, pipelined loops are non-reduction loops with no dependencies, giving $II=1$. However, if loop L10 in Listing 3 were pipelined, it would involve a reduction requiring $II\geq IL_{+}$, as the write of iteration $k$ must complete before the read of $k+1$.
(2) $\forall l\in\mathcal{L},loop_{l}\_pip\in\{0,1\}$
(3) $\forall l\in\mathcal{L},(1-loop_{l}\_pip)\times TC_{l,1}==1$
(4) $\forall s\in\mathcal{S},\sum_{l\in\mathcal{L}_{s}}loop_{l}\_pip\leq 1$
(5) $\forall s\in\mathcal{S},II_{s}=\sum_{l\in\mathcal{L}_{s}}loop_{l}\_pip\times II_{l}$
Coarse-grained unrolling  Coarse-grained unrolling applies only to non-reduction loops (Eq. 6). For instance, loop L3 in Listing LABEL:lst:distributed cannot be unrolled as it is a reduction loop. The unroll factor (UF) must divide the loop trip count and not exceed it (Eq. 7 and 8). However, as noted in Section 3, coarse-grained unrolling is skipped if the loop has only one statement (Eq. 9), as in Listing 1.
(6) $\forall s\in\mathcal{S}_{1},\forall l\in\mathcal{L}_{s}^{red},loop_{l}\_UF=1$
(7) $\forall l\in\mathcal{L},loop_{l}\_UF\%TC_{l,0}==0$
(8) $\forall l\in\mathcal{L},loop_{l}\_UF\leq TC_{l,0}$
(9) $\forall s\in\mathcal{S}_{1},\forall l\in\mathcal{L}_{s},loop_{l}\_UF=1$
On-chip memory  Constraints 10 and 11 ensure each array, represented by a boolean (Eq. 10), is cached at a single location per loop body, with only one boolean true per loop (Eq. 11). In Listing 3, array $A$ can be cached before loops (line 5) or after L6, L7, or L8 (lines 23, 25, 27). In Listing 2, $A$ is cached on-chip, matching the original array’s size (line 2). In Listing LABEL:lst:small_mem, $A$ is cached on-chip at line 18 with a smaller footprint. The on-chip footprint depends on the transfer location. For Listing LABEL:lst:small_mem, $A$ is transferred below loop L5 (trip count 48), so its footprint is $200\times\frac{240}{48}$.
(10) $\forall l\in\mathcal{L},\forall a\in\mathcal{A},cache_{l}\_arr_{a}\in\{0,1\}$
(11) $\forall a\in\mathcal{A},\forall s\in\mathcal{S},\sum_{l\in\mathcal{L}_{s}},cache_{l}\_arr_{a}=1$
(12) $\sum_{a\in\mathcal{A}}\sum_{l\in\mathcal{L}}cache_{l}\_arr_{a}\times ft\_array_{a}\_loop_{l}\leq Mem$
Array partitioning  Unrolling is applied only if all required data can be accessed in parallel, ensured by pragma array_partitioning. For each array $a$ and dimension $d$, the cyclic partitioning factor must exceed the unroll factor (Eq. 13) to ensure data is on separate banks. If the array is reused, partitioning must align with each unroll factor iterating through the dimension (Eq. 14). $AP_{a,d}$ represents the cyclic partitioning factor used in the code, with the total partitioning per array constrained by the user-defined limit (Eq. 15). For example, in Listing 2, $C$ is unrolled 200 times in the first dimension (lines 11, 12) and 4 times in the second (lines 13, 14) for the first loop body, and 200 times in the first dimension (lines 25, 26) for the second. This gives $AP_{C,1}=200$, $AP_{C,2}=4$, and $200\times 4=800\leq max_{part}$, with $max_{part}=1,024$.
(13) $\forall a\in\mathcal{A},\forall d\in\mathbb{N},\forall l\in C_{a_{d}},AP_{a,d}\geq TC_{l,2}$
(14) $\forall a\in\mathcal{A},\forall d\in\mathbb{N},\forall l\in C_{a_{d}},AP_{a,d}\%TC_{l,2}==0$
(15) $\forall a\in\mathcal{A},\prod_{d\in\mathbb{N}}AP_{a,d}\leq max_{part}$
DSP utilization  To manage DSP utilization, we use a user-defined flag, $reuse_{opt}$. If $reuse_{opt}=1$, optimistic reuse is assumed: identical operations reuse DSPs across loop bodies, with total usage given by the maximum DSPs per operation (Eq. 16). However, if the operations differ, as discussed in (Pouget et al., 2025, 2024), no reuse is assumed. If $reuse_{opt}=0$, no reuse is assumed, and DSPs are summed for each statement (Eq. 17). In both cases, total usage must not exceed available DSPs (Eq. 18 and 19). If a multiplication (*) uses 3 DSPs and an addition ($+$) uses 2 DSPs, in Listing 2, S0 needs $200\times 4\times 3$ DSPs for multiplications, while S1 requires $200\times 4\times 3+200\times 4\times 3+200\times 4\times 2$ DSPs for two multiplications and one addition. With optimistic reuse, DSP usage for multiplication is $\max(200\times 4\times 3,2\times 200\times 4\times 3)+200\times 4\times 2$, while pessimistic reuse results in $200\times 4\times 3+2\times(200\times 4\times 3)+200\times 4\times 2$.
(16) $DSPs\_used_{opt}=\sum_{op\in\{+,-,*,/\}}\max_{s\in\mathcal{S}}(DSP_{s_{op}}/II_{s})$
(17) $DSPs\_used_{pes}=\sum_{op\in\{+,-,*,/\}}\sum_{s\in\mathcal{S}}(DSP_{s_{op}}/II_{s})$
(18) $reuse_{opt}\times DSPs\_used_{opt}\leq DSP_{available}$
(19) $(1-reuse_{opt})\times DSPs\_used_{pes}\leq DSP_{available}$
4.3. Objective Function
We use the objective function similar to the approach described in (Pouget et al., 2025, 2024), tailoring it to meet the specific needs of our problem. Here, $Lat^{s}_{2}$ corresponds to the fine-grained unrolled level (level 2) for the statement $s$, $Lat^{s}_{1}$ denotes the pipeline tile that incorporates the fine-grained unrolled level (level 1), and recursively, $Lat_{0}$ encompasses $Lat_{1}$, enabling coarse-grained unrolling and facilitating on-chip data caching (level 0). The memory latency $Lat^{s}_{mem}$ for a statement $s$ corresponds to the time required to transfer the array from off-chip to on-chip and back for the arrays needed by a statement $s$. We assume the memory transfer is pipelined, allowing the transfer of one data element (with a maximum bitwidth of 512 bits) per cycle. Therefore, the latency at a given level is calculated as the array footprint divided by the burst size (bitwidth used for transferring the array), multiplied by the trip count of the loop that iterates over the function responsible for the transfer, based on the selected permutation of level 0. Additionally, if an array is fully transferred on-chip for a statement $s$ and reused by a subsequent statement $s{’}$ ’, no further transfers are needed. The burst size for an array must divide its last dimension. For instance, if we have an array $float\text{ }A[512][20]$, the maximum burst size for this array is $128$ bits. This is because $128$ is the largest power of $2$ that can evenly divide the size of the last dimension ($20\times 32$ bits).
$\left.\begin{cases}&Lat^{s}_{2}=IL_{par}+IL_{seq}\times\prod_{l\in\mathcal{L}^{red}}\times\log_{2}(TC_{l,2})\\ &Lat^{s}_{1}=Lat^{s}_{2}+II\times(TC_{l,1}-1)\\ &Lat^{s}_{0}=\prod_{l\in\mathcal{L}}\frac{TC_{l,0}}{loop_{l}\_UF}\times Lat^{s}_{1}\\ &L^{s}_{mem}=\sum_{l\in\mathcal{L}}\max_{a\in\mathcal{A}}(cache_{l}\_arr_{a}\times ft\_array_{a}\_loop_{l}\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\times TCs(perm_{s})/burst_{a})\\ &Lat_{s}=Lat^{s}_{0}+L^{s}_{mem}\end{cases}\right.$
Finally, the objective function $obj\_func=\sum_{s\in\mathcal{S}}Lat_{s}$ is defined as the sum of the latencies of each loop body.
5. Post-Optimization and Customization
The NLP determines the trip counts, array partitioning, and on-chip buffer sizes. Data transfer functions between off-chip and on-chip memory are automatically generated with the maximum burst size. Further optimizations are also applied to improve QoR. Optimization for Non-Constant Trip Count  To optimize loops with non-constant trip counts, we implement a code transformation that preserves the NLP’s estimation while simplifying compilation for the HLS compiler. We achieve this by replacing loops with non-constant trip counts with the maximal trip count computed using PoCC (PoCC, [n. d.]). Next, we replace all statements iterated by this loop with calls to a function. For example, instead of using $C[i][i]+=B[i][i]$, we replace it with $C[i][j]=f(C[i][j],B[i][i])$, where the function $f$ ensures compliance with the constraints of the non-constant trip count loop. This function returns the computed value if the constraints are satisfied; otherwise, it simply returns the original output value. This technique helps to reduce compilation time and achieve designs with a good QoR. Grouping Data Transfers  In the NLP, we already account for parallel transfers of arrays when they are transferred at the same level in the code (e.g., within the same loop). However, we do not consider parallel transfers between arrays belonging to different loop bodies. In our example, we first manage the transfer of arrays required for initializing the matrix, transferring the arrays for multiplication only after the initialization is complete. However, FPGA DRAM banks allow parallel transfers of different arrays. Consequently, we can optimize load and store operations during post-processing to improve memory transfer overlap. Code Transformation for HLS  To help the HLS compiler pipeline loops with the correct initiation interval (II), we simplify reductions at level 2. A variable accumulates the reduction within the pipeline, which is then added to the output. Additionally, we use the loop_flatten off pragma on all reduction loops above the pipeline. Although this adjustment could theoretically improve the QoR, in practice, flattening these loops can complicate the code and potentially increase the initiation interval (II) of the pipeline. Customization  Our framework is designed to be explainable and user-friendly, requiring the user to provide only the affine C/C++ code of the kernel to be optimized and the available resources. To fully benefit developers, our NLP formulation allows the template to be completely customized to the user’s needs. Each variable in the design space can be defined by the user if they have specific constraints. For example, in the gemm template (Listing 3), the user can set the permutations for level 0 while allowing the NLP to explore other parameters. Additionally, they can specify the on-chip size of matrix $A$ by directly defining a variable in the NLP, allowing for exploration within the defined constraints.
6. Evaluation
The goal of this evaluation is to demonstrate how effectively our approach applies code transformations and inserts hardware directives while respecting user-defined resource constraints. We compared our method against four other frameworks: AutoDSE (Sohrabizadeh et al., 2022b), NLP-DSE (Pouget et al., 2025), HARP (Sohrabizadeh et al., 2023), and ScaleHLS (Ye et al., 2022). Since generating a bitstream without significant manual intervention is challenging for some frameworks, we first demonstrate our ability to achieve a high-quality design by utilizing all available resources—ensuring fair comparison with the other frameworks—without considering place and route. We generate HLS reports using the vitis-flow, which accounts for memory transfers between off-chip and on-chip, necessitating the inclusion of memory transfer steps in ScaleHLS’s code. Thus, we show that our designs are synthesizable by targeting the resources of a single SLR, which eliminates the need for manual partitioning. Due to the time required for bitstream generation, we restrict our evaluation to a subset of the kernels.
6.1. Setup
We conducted experiments using kernels from Polybench/C 4.2.1 (PolyBench, [n. d.]), a CNN kernel, and matrix multiplication tasks similar to those in the BERT transformer model. All computations used single-precision floating-point data for direct comparison with frameworks like AutoDSE, NLP-DSE, and ScaleHLS, using medium-sized Polybench/C datasets. We also compare our approach with HARP (Sohrabizadeh et al., 2023), which enhances HLS design exploration using a Graph Neural Network (GNN) model to predict HLS tool behavior. HARP explores this space for one hour and synthesizes the top 10 designs. We use the medium-sized Polybench kernels found in their training set with double-precision floating-point data types, aiming to deploy HARP under optimal conditions. As detailed in (Pouget et al., 2025), HARP’s effectiveness is not universally applicable without the necessary fine-tuning or training, particularly when confronted with diverse problem sizes. We selected these problem sizes to highlight how our method performs in transforming code and inserting pragmas in cases where full unrolling is not feasible. We also demonstrate the effectiveness of our framework in tiling large tasks such as CNN and bert_3072_100, ensuring that only a small portion (a tile) of each array fits on-chip, in line with memory constraints. For the CNN kernel, the problem size and loop dimensions are the same as described in Section 2. For bert_n_m matrix multiplication, the dimensions are I = n, J = m, K = n (reduction). We used the AMD/Xilinx Vitis HLS compiler (Xilinx, 0232) to demonstrate our method’s effectiveness, generating reports with Vitis 2023.2. When enabling tree reduction, we selected the “funsafe math optimizations” option to allow commutative and associative reduction operators for logarithmic time reductions. For our hardware setup, we targeted the Xilinx Alveo U200 device operating at 250 MHz. We utilized the commercial solver Gurobi 11.0.0 (Gurobi Optimization, LLC, 2024) with AMPL.
6.2. Experimental Evaluation
To ensure a fair comparison, we modified kernels with non-constant trip counts for all benchmarks, as detailed in Section 5, except for ScaleHLS, which already handles such loops. We generated AutoDSE’s design space using the ds_generator command, applying all trip count factors for unroll and tile sizes. Since AutoDSE lacks array partitioning, its design space is slightly larger. AutoDSE was run with a 1,000-minute exploration timeout and a 180-minute HLS synthesis timeout. For NLP-DSE, we followed the original paper’s parameters (Pouget et al., 2025). HARP was run with the authors’ settings (Sohrabizadeh et al., 2023) using Vitis 2021.1 and tree reduction enabled. ScaleHLS assumes that all kernel data are already located in on-chip memory. To provide a fair comparison, we modified their code to handle off-chip to on-chip memory transfers. We optimized these transfers to 512 bits per cycle, with all arrays transferred in parallel. Each array was assigned to a different DRAM bank. However, due to the extensive manual modifications required, we did not extend this comparison to all kernels. For Sisyphus evaluation, we set the Gurobi solver timeout to 14,400 seconds per design. By default, we assume optimistic resource reuse between non-parallel statements. If DSP utilization exceeds limits, we rerun the NLP with pessimistic assumptions (no resource reuse) and regenerate the design. The maximum partitioning constraint, $max_{part}$, was set to 1024, per AMD/Xilinx guidelines. Each synthesis starts with a C-simulation to verify correctness. We generate Vitis reports using the vitis-flow, which counts memory transfers for all frameworks and considers the resources of the entire FPGA. For bitstream generation, we consider only one SLR to eliminate any manual intervention. We consider 60% of the resources of a single SLR to quickly identify a synthesizable design, minimizing the number of times AutoDSE needs to be run.
6.3. Comparison of the Methods
6.3.1. AutoDSE, NLP-DSE and Scale-HLS with tree reduction
Table 3 provides a comparison among AutoDSE (A-DSE), NLP-DSE (N-DSE), Scale-HLS (S-HLS) and Sisyphus (ours) with tree reduction. The T column indicates the specific transformations applied, with D representing distribution, T for data tiling, and P for permutation. The column Perf. Impr. of ours framework vs. displays the performance improvement of our framework compared to the three different frameworks.
GF/s Perf. Impr. of ours framework vs.
Kernel T Ours A-DSE N-DSE S-HLS
2mm D,P 175.49 431.17x 1.42x 5.50x
3mm D,P 140.83 80.75x 1.02x 3.86x
atax D 1.96 0.99x 1.00x 1.23x
bicg D,P 1.96 1.97x 1.97x 1.18x
doitgen D,P 54.05 1.35x 2.71x 8.45x
gemm D,P 203.39 1.84x 1.55x 5.81x
gemver P 17.62 4.58x 1.74x 9.39x
gesummv D 1.96 0.99x 1.00x 0.71x
mvt P 13.24 1.70x 1.70x 1.26x
symm D,P 240.50 8.28x 8.28x 2,704.51x
syr2k D,P 254.69 5.52x 1.86x 174.47x
syrk D,P 161.20 6.55x 2.32x 358.95x
trmm D,P 148.53 8,345.10x 8,312.42x 2,376.91x
Average 108.88 683.91x 641.46x 434.78x
Geo Mean 40.13 9.17x 3.47x 16.15x
Table 3. Throughput (GF/s) Comparison with Tree Reduction Using Vitis HLS
Across all evaluated kernels, we consistently achieve performance that is comparable to or better than alternatives, except for the atax, and gesummv kernels, where we observe a slowdown of less than 5%. For 3mm, symm and trmm we had to apply the constraint indicating no reuse (Eq. 17), as the constraint with optimistic reuse (Eq. 16) resulted in kernels with resource over-utilization. For Gesummv, ScaleHLS is 1.41x faster. However, 91% of our design’s execution time is spent on data transfers. ScaleHLS benefits from 512-bit data transfers, which we manually applied for testing. Our framework uses 64-bit chunks, the largest bitwidth that divides the array. With 512-bit transfers, our framework could achieve 14.50 GF/s. For other memory-bound kernels like Atax and Gesummv, the slight performance differences observed with other frameworks are largely due to how Merlin handles memory transfers. Sisyphus averages a speedup of 683.91x, 641.46x and 434.78x over AutoDSE, NLP-DSE and Scale-HLS respectively, across the evaluated kernels. In terms of geometric mean, Sisyphus achieves a speedup of 9.17x, 3.47x and 16.15x. Preserve the original schedule when necessary  For atax, Sisyphus preserves the original schedule and inserts pragmas, effectively maintaining the existing schedule while efficiently incorporating pragmas, similar to two DSEs focused on pragma insertion. Code transformation for memory-bound kernel  For mvt and gemver, both memory-bound kernels, effective optimization entails loop permutations to prevent redundant array transfers and accommodate array partitioning within the constraints of the AMD/Xilinx compiler’s 1024 limit. Sisyphus adeptly manages these loop permutations and pragma insertions, guaranteeing optimized performance for these kernels. For bicg, fully distributing and arranging the loops appropriately can enhance parallelization capabilities. Allow to achieve a parallelism not achievable without loop transformation  For doitgen, symm, syrk, syr2k, and trmm, the original schedule imposes constraints on the achievable level of parallelism. This limitation highlights the performance enhancements made possible by Sisyphus. We observe QoR improvements for these kernels because the code transformations enable a combination of loop unrolling that cannot be achieved even with maximal loop distribution, as each loop can be partially unrolled.
6.3.2. AutoDSE and NLP-DSE without tree reduction
Transformations without tree reduction, which may be needed to preserve floating-point data precision, must be different since the optimizations differ when tree reductions are not involved. Table 4 provides a comparison among AutoDSE (A-DSE), NLP-DSE (N-DSE) and Sisyphus (ours) without tree reduction.
GF/s Perf. Impr. Ours vs.
Kernel T Ours A-DSE N-DSE
2mm D,P 171.07 817.42x 1.44x
3mm D,P 140.11 511.64x 1.04x
atax D 1.86 0.98x 1.00x
bicg D,P 1.86 1.92x 1.93x
doitgen D,P 54.05 5.46x 2.71x
gemm D,P 210.61 9.25x 4.01x
gemver P 15.01 37.28x 8.73x
gesummv D,P 1.81 0.97x 0.97x
mvt P 9.75 1.42x 1.42x
symm D,P 240.55 8.28x 7.32x
syr2k D,P 429.34 17.61x 3.14x
syrk D,P 333.03 13.54x 4.78x
trmm D,P 130.90 7,362.68x 7,333.85x
Average 133.84 676.03x 567.10x
Geo Mean 41.62 18.50x 4.49x
Table 4. Throughput (GF/s) Comparison without Tree Reduction Using Vitis HLS
Sisyphus achieves an average speedup of 676.03x over AutoDSE and 567.10x over NLP-DSE across kernels without tree reduction, with geometric means of 18.50x and 4.49x, respectively. Performance improvement is higher without tree reduction, especially for gemm (Listing 2) and gemver, as optimizing reduction loops becomes harder, increasing the need for code transformations. The difference between syrk and syr2k with tree reduction occurs because the HLS compiler achieves II=2 instead of the expected II=1, while without tree reduction, II=1 is achieved.
6.3.3. AutoDSE and NLP-DSE for Bert and CNN
Table 5 compares AutoDSE (A-DSE), NLP-DSE (N-DSE), and Sisyphus (ours). The TR column shows the tree reduction status. For bert_100_64, we observed a minor decrease in QoR, attributed to the NLP selecting a configuration where the pipeline was not applied as expected.
GF/s Perf. Impr. Ours vs.
Kernel TR T Ours A-DSE N-DSE
bert_100_64 1 D,P 81.22 0.99x 0.95x
bert_100_64 0 D,P 76.76 1.09x 1.09x
bert_100_768 1 D,P 218.08 1.12x 1.19x
bert_100_768 0 D,P 216.95 1.13x 1.13x
bert_100_3072 1 D,P 241.59 1.21x 1.21x
bert_100_3072 0 D,P 241.06 1.22x 1.22x
bert_3072_100 1 D,P,T 314.89 21.85x 24.29x
bert_3072_100 0 D,P,T 82.41 989.65x 1,156.02x
cnn 1 D,P,T 314.15 7.38x 8.23x
cnn 0 D,P,T 341.20 8.09x 10.73x
Average 212.83 103.37x 120.61x
Geo Mean 185.32 4.38x 4.69x
Table 5. Throughput (GF/s) Comparison Using Vitis HLS
Code transformation with tiling  The selection of loop order and tile size is critical for determining which array sizes, particularly for large problem size like bert_3072_100 and CNN layers, should be cached on-chip and where they should be transferred. In these instances, the NLP aims to enhance data reuse and minimize unnecessary memory transfers while identifying the theoretical optimal balance with parallelism.
6.3.4. HARP
In Table 6 we compare the throughput and resource utilization (BRAM, DSP, LUT, FF) achieved with HARP (Sohrabizadeh et al., 2023).
GF/s Resource Utilization (%) Perf. Impr.
BRAM,DSP,LUT,FF
Kernel HARP Ours HARP Sisyphus
atax 1.72 1.96 78,52,49,50 38,65,53,33 1.14x
bicg 0.92 1.96 75,25,35,36 38,65,48,31 2.13x
gemm 125.59 87.86 29,80,55,40 32,48,36,23 0.70x
gemver 1.66 9.42 29,28,35,19 34,8,31,11 5.69x
mvt 7.07 8.29 40,78,43,30 53,44,57,30 1.17x
Average 27.39 20.71 1.45x
Geo Mean 4.71 6.28 1.33x
Table 6. Comparison with HARP Using Vitis HLS
The enhanced performance seen in both bicg and gemver can be attributed to the same underlying factor observed in the comparison with AutoDSE and NLP-DSE. Regarding mvt, HARP produced comparable performance to our own due to their use of the Merlin compiler, which facilitates loop permutations when partially unrolling consecutive loops. This allows them to find the same schedule that our framework identified. However, for gemm, the design discovered by HARP falls outside our specified parameter space, due to our restriction on maximum array partitioning.
6.4. Single SLR Onboard Evaluation
GF/s RU: BRAM,DSP,LUT(K),FF(K) Perf. Impr.
Kernel Ours Ours AutoDSE
2mm 30.57 510,556,213,276 353.5,963,287,292 76.98x
3mm 29.89 611,984,230,300 470,1117,278,306 72.34x
atax 1.03 450,173,240,250 630.5,452,170,212 4.63x
bicg 1.02 451,173,238,265 867.5,196,168,217 1.80x
Table 7. Comparison of Onboard Evaluation for 1 SLR
The resource utilization (RU) is detailed in the thousands for both LUTs and FFs. URAM utilization is excluded as no kernels use it. And the last column show the performance improvement we achieve with Sisyphus over AutoDSE. For the kernels evaluated on board, we can see that our generated design is both faster and consumes fewer resources compared to AutoDSE.
6.5. Latency Estimation and Scalability
Implementing the post-optimization strategy from Section 5 results in a loss of the lower bound discussed in (Pouget et al., 2024, 2025), mainly because we exclude communication overlap modeling between loop bodies in the NLP to avoid significantly increasing search time. We also do not incorporate the loop_flatten pragma, even if the compiler may automatically apply it for loops without reductions, as we deactivate it for reduction loops. Although this leads to a lower bound loss, we prefer not to depend on an optimization that may not be consistently implemented. As noted in Section 5, our post-optimization focuses primarily on memory transfer overlap, so our final latency values do not directly reflect the model’s initial predictions. However, our model achieves (when predicting the latency of the final design generated) an average prediction error of 8% across evaluated designs, with a 3% error in geometric mean. For certain designs, like syrk and syr2k, where the compiler achieves an II=2 rather than the expected II=1, prediction errors are higher (49% and 32%, respectively) due to an optimistic estimation of II. The average solving time of the NLP was 455 seconds, while the geometric mean stood at 6.7 seconds. Only the kernel 3mm with tree reduction timeout at 4h. 32/36 NLP’s problem are solve in less than 1 minutes. It is important to note that the solver provides a feasible solution early in the process, though it may not be theoretically optimal. As the search continues, the solution improves, and once the solver completes, the solution is theoretically optimal. Therefore, it is possible to set a timeout to obtain a feasible, but not optimal, solution more quickly.
7. Related Work
Various code transformations have been devised for CPUs (Bondhugula et al., 2008), GPUs (Verdoolaege et al., 2013), FPGAs (Pouchet et al., 2013b; Zhao et al., 2022; Zhao and Cheng, 2021; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) or for all (Zheng et al., 2020). While code transformations for CPUs and GPUs yield remarkable results tailored to their respective architectures, they may not be inherently suitable for FPGA targets aimed at achieving high parallelism. Regarding (Zhao et al., 2022; Zhao and Cheng, 2021), they employ Pluto on different scopes of the kernel for code transformation, yet its pragma insertion capabilities are limited, making it incomparable to our work. Conversely, (Pouchet et al., 2013b; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) pursue a distinct objective from ours. (Pouchet et al., 2013b) aims to minimize communication between off-chip and on-chip, which results in better QoR than ours for memory-bound kernels. The works (Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) concentrate on code transformations aimed at enhancing pipelining techniques. However, these objectives may not be suitable for computation-bound kernels requiring high levels of parallelization. Ansor (Zheng et al., 2020) achieves impressive results for CPU and GPU architectures; however, there are notable differences with our approach. Ansor requires multiple evaluations to perform its design-space exploration (DSE) and update its cost model, which accelerates the DSE process but is not scalable for FPGA, where a single evaluation can take minutes or even hours. Furthermore, Ansor’s approach to code transformations and pragma insertions (which are more tailored for CPU/GPU) addresses these problems separately, potentially resulting in suboptimal designs. The choice of tile sizes significantly impacts the QoR. In line with our methodology, (Liu et al., 2017a; Li et al., 2021) employ a cost model to determine the tile size. However, while (Liu et al., 2017a) focuses on minimizing communication overhead, our approach differs. On the other hand, (Li et al., 2021) investigates tiles size selection for Convolutional Neural Networks (CNNs) with three-level CPU caching.
8. Limitations
Sisyphus is designed to optimize affine loop nests. For applications with non-affine code, we can partition the code into affine and non-affine sections. Sisyphus handles affine regions effectively, while tools like AutoDSE (Sohrabizadeh et al., 2022b) or HARP (Sohrabizadeh et al., 2023) can optimize non-affine parts. The optimized sections are then integrated using a dataflow model, either with FIFOs for data exchange or by employing TAPA (Guo et al., 2023), which facilitates dataflow programming. This combined approach ensures that both affine and non-affine portions of the application are handled in an optimal manner, achieving a balance between different types of optimizations and seamlessly integrating them for improved overall performance.
9. Conclusion
In this paper, we introduced Sisyphus, a unified framework that streamlines hardware design through HLS. By integrating code transformation, pragma insertion, and tile size selection into a single optimization, Sisyphus addresses challenges in existing HLS workflows, which often rely on manual and disjointed methods. Sisyphus defines a design space in the form of a template. Using NLP, it efficiently explores this space, quickly identifying transformations and pragma configurations to achieve designs with good QoR. Our evaluation demonstrates that Sisyphus outperforms AutoDSE, NLP-DSE, and ScaleHLS in terms of quality of result. Acknowledgments -   This work was supported by the NSF award #CCF-2211557. It is also supported by CDSC industrial partners and the AMD [^1] J. Cong has a financial interest in AMD.
/HACC Program.